OpenStudy (anonymous):

One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together? 4 min 5 min 6 min

OpenStudy (anonymous):

I think it has to be 6. If they worked on the same rate, it should drop by a half. As one is slower than the other, we should have a bit more than a half, i.e., 6 min.

hero (hero):

Hint: \[\frac{P1 \times P2}{P1 + P2}\]

hero (hero):

@bmp good job giving the answer.

hero (hero):

@myininaya @TuringTest

hero (hero):

@TuringTest

OpenStudy (anonymous):

first person does 1/10 of the job in a minute. Second person does 1/15 of a job in a minute. Now you need to add them together to see what they can do together so (1/10) + (1/15) = 1/6 Now I let "t" stand for time together so we are looking for 1/t to see what together they can do per minute. 1/6=1/t so t = 6 mins in case you don't get what @Hero said. sorry I was already typing and figured I'd post anyway lol.

hero (hero):

P1 = Person 1 P2 = Person 2 If it wasn't explicitly obvious, it's a formula that helps find the answer.

OpenStudy (anonymous):

Sorry for breaking the CoC. I generally never give away the answer but as I was unsure, I didn't think there was a problem. Anyway, I understand if I get a warning or something else by this. Sorry again.